# Shock waves in nonlinear transmission lines

### v1

@@ II The circuit equations
$Q$ $=Q(V),$ (4a)
$\Phi$ $=\Phi(I).$ (4b)
Further on we’ll consider $Q$ and $\Phi$ as known functions of $V$ and $I$ respectively. Note however, that sometimes it will be convenient for us to invert the relations (4) and consider $V$ and $I$ as known functions of $Q$ and $\Phi$ respectively. We’ll switch between the representations without further notice.
@@ III The ”sound” and the shocks
$C_{0}$ $=\left.\frac{dQ}{dV}\right|_{V=V_{0}},$ (7a)
$L_{0}$ $=\left.\frac{d\Phi}{dI}\right|_{I=I_{0}}.$ (7b)
@@
The second type of solutions, we are interested in, is shock waves, satisfying the conditions (see Section V)
@@
where the indices 1 and 2 refer to the quantities before and after the shock respectively, and $U_{21}$ is the shock wave speed. Equations (11) mean, in particular, that from the five parameters ($V_{1},V_{2},I_{1},I_{2},U$) - three are the free parameters, and the other two are defined by the equations.
The indices 1 and 2 enter in a symmetrical way in the system 11, so Eqs. 11 as they stand can describe propagation of the shock in both directions. However, in Section V) we’ll show that the shock always propagates in such a way that whitham ; kogan4 ; kogan5
@@ IV.1 Nonlinear transmission line
where (in) stands for the incident ”sound” wave and (r) for the reflected ”sound” wave. We also have to take into account the change of the shock speed due to the ”sound” waves
@@
(the difference in the signs is because of the opposite directions of propagation of the two waves) and dividing the equations (19) to each other we obtain
$Z_{ab}=Z_{a}\frac{u_{in}i^{(in)}+\left(u_{a}+U_{ab}\right)i^{(r)}}{u_{r}i^{(in)}-\left(u_{a}+U_{ab}\right)i^{(r)}},$ (21)
where $u_{in}=u_{a}-U$ is the speed of the incident ”sound” wave relative to the shock wave, $u_{r}=u_{a}+U$ is the speed of the reflected ”sound” wave relative to the shock wave and we introduced the impedance of the shock wave
$Z_{ab}\equiv\frac{V_{a}-V_{b}}{I_{a}-I_{b}}=\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (22)
Solving (21) we obtain for the reflection coefficient
$R\equiv\frac{i^{(r)}}{i^{(in)}}=\frac{u_{in}\left(Z_{ab}-Z_{a}\right)}{u_{r}\left(Z_{ab}+Z_{a}\right)}.$ (23)
As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $\Phi_{a}\to\Phi_{b},V_{a}\to V_{b}$, in other words, when the shock wave itself nearly becomes the ”sound” wave.
@@
Taking into account the relations
@@
and dividing the equations (29) to each other we obtain
$Z_{ab}=\frac{u_{a}u_{in}Z_{b}i^{(in)}-u_{b}u_{t}Z_{a}i^{(t)}}{-u_{a}u_{in}i^{(in)}+u_{b}u_{t}i^{(t)}},$ (31)
where $u_{in}=u_{b}+U$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U$ is the speed of the transmitted ”sound” wave relative to the shock wave. Solving (31) we obtain for the transmission coefficient
$T\equiv\frac{i^{(t)}}{i^{(in)}}=\frac{u_{a}u_{in}\left(Z_{ab}+Z_{b}\right)}{u_{b}u_{t}\left(Z_{ab}+Z_{a}\right)}.$ (32)
As one could have expected, the transmission coefficient goes to 1 when the intensity of the shock wave decreases, that is when $\Phi_{a}\to\Phi_{b},V_{a}\to V_{b}$, in other words, when the shock wave itself nearly becomes the ”sound” wave.
Note that Eqs. (21) and (31) can be written down respectively in the form
$u_{in}\left(Z_{ab}-Z_{a}\right)i^{(in)}$ $=u_{r}\left(Z_{ab}+Z_{a}\right)i^{(r)},$ (33a)
$u_{a}u_{in}\left(Z_{ab}+Z_{b}\right)i^{(in)}$ $=u_{b}u_{t}\left(Z_{ab}+Z_{a}\right)i^{(t)}.$ (33b)
We wonder whether these equations could have been written down at an earlier stage by considering propagation of the waves in the frame of reference moving with the shock wave.
IV.2 Half-nonlinear transmission line
IV.2.1 Linear capacitor
@@ V.1 Nonlinear transmission line
The point $(Q_{1},\Phi_{1})$ is the unstable fixed point of the system (48). This means
@@
On the other hand, the point $(Q_{2},\Phi_{2})$ is the stable fixed point, hence we obtain
@@
V.2 Half-nonlinear transmission line
Further analysis of the fully nonlinear line we relegate to Appendix A) and here consider transmission line with the linear capacitor. Because of the symmetry of Eq. (48) given by (45) (to which we should add $1/R_{L}\leftrightarrows R_{C}$), the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section..
@@ V.2 Half-nonlinear transmission line
$\frac{LR_{C}}{R_{L}}\frac{d^{2}Q}{d\tau^{2}}+R_{C}\frac{dQ}{d\tau}+\frac{L}{R_{L}}\frac{dV}{d\tau}=F,$ (59)
@@
$F\equiv\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q-V+\frac{V_{1}Q_{2}-V_{2}Q_{1}}{Q_{2}-Q_{1}}.$ (60)
Equation (59) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well defined by (60). The boundary conditions (49a) tell that the particle starts its motion at $\tau=-\infty$ at the potential maximum at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$, thus it is crucial to keep the nonlinear dependence of $Q$ upon $V$ in 60. We may present these maximum/minimum conditions as
@@
Equation (59) is simplified (in a trivial way) when there is a leading mechanism of dissipation, that is when either $R_{C}Q\ll LV/R_{L}$ or $R_{C}Q\gg LV/R_{L}$. In each of these cases, in the l.h.s. of (59) a single term with the first derivative should be left, and the equation can be easily integrated in quadratures (in some cases in elementary functions). Thus we obtained Eqs. (48a) and (48b) of our paper kogan3 .
V.2.1 Weak shocks
Alternatively, Eq. (59) is simplified (in a less trivial way) for weak shock, when $|V_{1}-V_{2}|$ is small. We can put in the l.h.s. of the equation
$\frac{dQ}{d\tau}=C_{s}\frac{dV}{d\tau},$ (62)
where $C_{s}$ is a differential capacitance assumed to be constant for the voltages considered, and (59) becomes
$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=F(V),$ (63)
where $\tau_{R}\equiv R_{C}C_{s}+L/R_{L}$ and $T^{2}\equiv LC_{s}R_{C}/R_{L}$. The solution of (63) should satisfy the boundary conditions (49a), rewritten as
$\lim_{\tau\to-\infty}V(\tau)=V_{1};\hskip 14.22636pt\lim_{\tau\to+\infty}V(\tau)=V_{2}.$ (64)
@@ V.2.1 Weak shocks
$F(V)=\frac{(V-V_{1})\left[Q(V)-Q_{2}\right]-\left\{2\leftrightarrows 1\right\}}{Q_{2}-Q_{1}},$ (65)
where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged. For weak shocks we can approximate $Q(V)$ in (65) as
$Q(V)=Q_{0}+C_{0}(V+\beta V^{2}),$ (66)
and (63) becomes
$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=\beta(V-V_{1})(V-V_{2}).$ (67)
In the Appendix B we’ll show that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{25\beta(V_{1}-V_{2})}{6},$ (68)
@@
$V=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi\tau\right)+1\right]^{2}},$ (69)
@@
Approximation (66) can be inadequate, if $V_{1},V_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. So we are motivated to generalize (66) to
$Q=Q_{0}+C_{0}(V+\beta V^{2}+\gamma V^{3}).$ (71)
@@
$T^{2}\frac{d^{2}V}{d\tau^{2}}+\tau_{R}\frac{dV}{d\tau}=\gamma(V-V_{1})(V-V_{2})(V+V_{3}),$ (72)
where $V_{3}=\beta/\gamma+V_{1}+V_{2}$.
In the Appendix B we’ll show that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(V_{1}+V_{2}+2V_{3}\right)^{2}}{2},$ (73)
@@
$V=V_{2}+\frac{V_{1}-V_{2}}{\exp\left(\psi\tau\right)+1},$ (74)
@@
$\psi=\frac{\tau_{R}\left(V_{1}-V_{2}\right)}{T^{2}\left(V_{1}+V_{2}+2V_{3}\right)}.$ (75)
@@ Acknowledgements.
We are grateful to A. M. Kamchatnov for sending us the manuscript of his (to be published) excellent book ”Theory of Nonlinear Waves”, which motivated the continuation of our previous studies. We are very grateful to J. Cuevas-Maraver, M. Goldstein, A. Kamchatnov, B. Malomed, A. Novic-Cohen and V. Verbuk for the insightful discussions. We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research. The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems, ICTP, Trieste. The Author is grateful to the Center for the hospitality.
Appendix A Shock profile for the fully nonlinear transmission line
Taking into account the boundary conditions we can present (48) as
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=UQ-I(\Phi)-UQ_{1}+I_{1},$ (76a)
$R_{C}\frac{dQ}{d\tau}$ $=U\Phi-V(Q)-U\Phi_{1}+V_{1},$ (76b)
Equation (76) allows us to study the profile of the shock wave in the general case.
$V(Q)=U\Phi-U\Phi_{1}+V_{1}.$ (77)
Solving (77) with respect to $Q$ and substituting into (76a) we obtain the equation which can be easily integrated in quadratures.
In the second case we equate the r.h.s. of (76a) to zero and obtain
$I(\Phi)=UQ-UQ_{1}+I_{1}.$ (78)
Solving (78) with respect to $\Phi$ and substituting into (76b) we obtain the equation which can be integrated in quadratures.
If both $R_{C}>0$ and $R_{L}<\infty$, from (76) we may obtain the equation
$\frac{1}{R_{L}R_{C}}\frac{d\Phi}{dQ}=\frac{UQ-I(\Phi)-UQ_{1}+I_{1}}{U\Phi-V(Q)-U\Phi_{1}+V_{1}}.$ (79)
If we can solve (79), we can substitute the solution into one of the equations (76) and integrate the resulting equations in quadratures to obtain the profile of the shock wave.
Appendix B Integration of the ODE which don’t contain explicitly the independent variable in terms of elementary functions
Looking at Eq. (63) we realize that it doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $V$ as the new independent variable and $E\equiv dV/d\tau$ as the new dependent variable. In the new variables (63) takes the form
$T^{2}E\frac{dE}{dV}+\tau_{R}E=F(V).$ (80)
@@ Appendix A Shock profile for the fully nonlinear transmission line
$\lim_{V\to V_{1},V_{2}}E(V)=0.$ (81)
Note that Eq. (80) and the boundary conditions (81) are symmetric with respect to interchange of the indexes $2$ and $1$, while Eq. (63) and the boundary conditions (49a) - certainly not. So to the boundary conditions (64) we should add the demand (61) (or opposite inequality for $V=V_{1}$). Substituting (61) into (80) we obtain the demand written in the new variables as
$\left.\frac{T^{2}}{2}\frac{d^{2}E}{dV^{2}}\right|_{V=V_{2}}+\left.\tau_{R}\frac{dE}{dV}\right|_{V=V_{2}}<0.$ (82)
@@
$T^{2}E\frac{dE}{dV}+\tau_{R}E=\gamma(V-V_{1})(V-V_{2})(V+V_{3}).$ (83)
The boundary conditions (81) and the r.h.s. of (83) prompt the idea to look for $E$ as the polynomial function proportional to $\left(V-V_{1}\right)\left(V-V_{2}\right)$. Because the r.h.s. of (83) is the third order polynomial of $V$, the multiplying factor should be a constant. That is the natural inspired guess is kogan4 ; kogan5
$E(V)=\frac{\psi}{V_{1}-V_{2}}\left(V-V_{1}\right)\left(V-V_{2}\right)$ (84)
@@
$\frac{T^{2}\psi^{2}}{\left(V_{1}-V_{2}\right)^{2}}\left(2V-V_{1}-V_{2}\right)+\frac{\tau_{R}\psi}{V_{1}-V_{2}}=\gamma\left(V+V_{3}\right).$
@@
Hence the condition on the parameters of the problem, for (84) to be the solution of (83) is given by (73), and the constant multiplier in (84) should is given by (75). Substituting (84) into the equation
$\frac{dV}{d\tau}=E(V)$ (86)
we obtain the solution (74).
For $Q(V)$ given by (66) we obtain, instead of (83), Abel equation of the second kind polyanin
$T^{2}E\frac{dE}{dV}+\tau_{R}E=\beta(V-V_{1})(V-V_{2}).$ (87)
@@
Counting the powers of $V$ in the r.h.s. of (87) we come to the conclusion that if $E$ is a polynomial function of $V$, it should be a 3/2 order polynomial function. And the boundary conditions (81) tell us that this polynomial should zero out at $V=V_{1,2}$. To acquire an inspired guess how the polynomial may look like, let us return to the solution of (83) we obtained previously and split in two one multiplier in the solution
$\frac{V-V_{1}}{V_{1}-V_{2}}=\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}-1\right]\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}+1\right].$
All that was said above being taken together, prompts us to check up the solution of (87) in the form kogan4 ; kogan5
$E(V)=2\chi\left[\left(\frac{V-V_{2}}{V_{1}-V_{2}}\right)^{1/2}-1\right]\left(V-V_{2}\right).$ (89)
@@
Substituting the ansatz into (87) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (68) and $\chi$ is given by (70). Substituting (89) into (86) we obtain (69)
Actually, because of the symmetry of Eq. (80) and the boundary conditions (81) with respect to interchange of the indexes $2$ and $1$, instead of (89) and the condition (68) we could as well have obtained those with the indexes $2$ and $1$ interchanged. However, one can easily see that (89) satisfies the demand (82), and (89) with the indexes 1 and 2 interchanged - doesn’t.

### v3

@@ II The circuit equations
$V$ $=V(Q),$ (4a)
$I$ $=\Phi(\Phi).$ (4b)
Further on we’ll consider $V$ and $I$ as known functions of $Q$ and $\Phi$ respectively.
@@ III The ”sound” and the shocks
$C_{0}$ $=\left.\frac{dQ}{dV}\right|_{Q=Q_{0}},$ (7a)
$L_{0}$ $=\left.\frac{d\Phi}{dI}\right|_{\Phi=\Phi_{0}}.$ (7b)
@@
The second type of solutions, we are interested in, is shock waves, containing (in the approximation of the present Section) discontinuities, satisfying the conditions (see Section V)
@@
where
$V_{a,b}\equiv V(Q_{a,b}),\hskip 28.45274ptI_{a,b}\equiv I(\Phi_{a,b}),$ (12)
and $U_{21}$ is the shock wave speed. The indices 1 and 2, which refer to the quantities before and after the shock respectively, enter in the system (11) in a symmetrical way. However, in Section V we’ll show that the shock always propagates in such a way that whitham ; kogan4 ; kogan5
@@
The shock is supersonic with respect to the phase before the front, but subsonic with respect to the phase behind the front.
@@ IV The ”sound” reflection and transmission
where (in) stands for the incident and (r) - for the reflected ”sound” wave. We also have to take into account the change of the shock speed due to the ”sound” waves
@@
Dividing the equations (20) to each other we obtain
$u_{in}v^{(in)}-u_{r}v^{(r)}=Z_{ab}\left(u_{in}i^{(in)}-u_{r}i^{(r)}\right),$ (21)
where $u_{in}=u_{a}-U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, $u_{r}=u_{a}+U_{ab}$ is the speed of the reflected ”sound” wave relative to the shock wave and we introduced the impedance of the shock wave
$Z_{ab}\equiv\frac{V_{a}-V_{b}}{I_{a}-I_{b}}=\frac{\Phi_{a}-\Phi_{b}}{Q_{a}-Q_{b}}.$ (22)
@@
(the difference in the signs is because of the opposite directions of propagation of the two waves) and solving (21) we obtain for the reflection coefficient
$R\equiv\frac{i^{(r)}}{i^{(in)}}=\frac{\left(Z_{ab}-Z_{a}\right)u_{in}}{\left(Z_{ab}+Z_{a}\right)u_{r}}.$ (24)
@@
Dividing the equations (30) to each other we obtain
$u_{a}u_{in}v^{(in)}-u_{b}u_{t}v^{(t)}=Z_{ab}\left(u_{a}u_{in}i^{(in)}-u_{b}u_{t}i^{(t)}\right),$
where $u_{in}=u_{b}+U_{ab}$ is the speed of the incident ”sound” wave relative to the shock wave, and $u_{t}=u_{a}+U_{ab}$ is the speed of the transmitted ”sound” wave relative to the shock wave. Taking into account the relations
@@
and solving (IV) we obtain for the transmission coefficient
$T\equiv\frac{i^{(t)}}{i^{(in)}}=\frac{\left(Z_{ab}+Z_{b}\right)u_{a}u_{in}}{\left(Z_{ab}+Z_{a}\right)u_{b}u_{t}}.$ (33)
As one could have expected, the modulus of the ”sound” reflection coefficient is less than one, and it goes to zero when the intensity of the shock wave decreases, that is when $Q_{a}\to Q_{b},\Phi_{a}\to\Phi_{b}$, in other words, when the shock wave itself is nearly the ”sound” wave. Similarly, in these circumstances the transmission coefficient goes to 1.
Looking back at (21) and (IV) we wonder whether these equations could have been written down by considering propagation of the waves in the frame of reference moving with the shock wave.
IV.1 Half-nonlinear transmission line
IV.1.1 Linear capacitor
@@ V Lossy transmission line
From the five parameters ($Q_{1},Q_{2},\Phi_{1},\Phi_{2},U$) - three are arbitrary constants, and the other two are defined by the equations.
@@
$(Q_{1},\Phi_{1})$ is the semistable fixed point of the system (48), and $(Q_{2},\Phi_{2})$ is the stable fixed point. (Other options would have demanded fine tuning of the parameters.) This means
@@
Similarly for the point $(Q_{2},\Phi_{2})$ we obtain
@@
Note that the kinks, which we considered in our previous publications in addition to shocks but do not consider here, correspond to both fixed points being semistable. Thus the kinks turn out to be supersonic both with respect to the phase before the front and the phase behind the front.
V.1 Half-nonlinear transmission line
Further analysis of the fully nonlinear line we relegate to Appendix B) and here consider transmission line with the linear capacitor. Because of the symmetry of Eq. (48) given by (45) (to which we should add $1/R_{L}\leftrightarrows R_{C}$), the case of the transmission line with linear capacitors and nonlinear inductors (JTL first and foremost) would be identical to the one considered in this Section..
@@ V.1 Half-nonlinear transmission line
$\frac{LR_{C}}{R_{L}}\frac{d^{2}Q}{d\tau^{2}}+R_{C}\frac{dQ}{d\tau}+\frac{L}{R_{L}}\frac{dV(Q)}{d\tau}=F(Q),$ (59)
@@
$F(Q)\equiv\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q-V(Q)+\frac{V_{1}Q_{2}-V_{2}Q_{1}}{Q_{2}-Q_{1}}.$ (60)
To (59) we should add the boundary conditions (49a). (The third arbitrary constant appears when we connect $\Phi$ with the found $Q$.)
Equation (59) describes motion with friction of a fictitious Newtonian particle, $Q$ being the coordinate of the particle, in the potential well defined by (60). The boundary conditions (49a) tell that the particle starts its motion at $\tau=-\infty$ at the potential maximum at $Q=Q_{1}$ and ends - at $\tau=+\infty$ at the potential minimum at $Q=Q_{2}$. We may present these maximum/minimum conditions as
@@
V.1.1 Weak shocks
@@
$F(Q)=\frac{\left[V(Q)-V_{1}\right]\left(Q-Q_{2}\right)-\left\{2\leftrightarrows 1\right\}}{Q_{2}-Q_{1}},$ (62)
where the notation $\left\{2\leftrightarrows 1\right\}$ means the preceding term with the indexes $2$ and $1$ interchanged. For weak shocks, when $|Q_{1}-Q_{2}|$ is small, we can approximate $V(Q)$ for $Q$ between $Q_{1}$ and $Q_{2}$ as
$V(Q)=\text{const}+\frac{Q+\beta Q^{2}}{C_{s}},$ (63)
where $\beta$ and $C_{s}$ are assumed to be constant. We also may take $V(Q)$ in the last term in the l.h.s. of (59) in the linear approximation. So finally Eq. (59) becomes
$T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\beta(Q-Q_{1})(Q-Q_{2}),$ (64)
where
$\tau_{R}$ $\equiv R_{C}C_{s}+\frac{L}{R_{L}},$ (65a)
$T^{2}$ $\equiv\frac{LC_{s}R_{C}}{R_{L}}.$ (65b)
One can check up that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{25\beta(Q_{1}-Q_{2})}{6},$ (66)
@@
$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\left[\exp\left(\chi\tau\right)+1\right]^{2}},$ (67)
@@
Approximation (63) can be inadequate, if $Q_{1},Q_{2}$ are close to zero and, as it normally happens, the charge is an odd function of the voltage. So we are motivated to generalize (63) to
$V(Q)=\text{const}+\frac{Q+\beta Q^{2}+\gamma Q^{3}}{C_{s}}.$ (69)
@@
$T^{2}\frac{d^{2}Q}{d\tau^{2}}+\tau_{R}\frac{dQ}{d\tau}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}),$ (70)
where $Q_{3}=\beta/\gamma+Q_{1}+Q_{2}$. One can check up that when the parameters of the problem satisfy the condition
$\frac{\tau_{R}^{2}}{T^{2}}=\frac{\gamma\left(Q_{1}+Q_{2}+2Q_{3}\right)^{2}}{2},$ (71)
@@
$Q=Q_{2}+\frac{Q_{1}-Q_{2}}{\exp\left(\psi\tau\right)+1},$ (72)
@@
$\psi=\frac{\tau_{R}\left(Q_{1}-Q_{2}\right)}{T^{2}\left(Q_{1}+Q_{2}+2Q_{3}\right)}.$ (73)
In the Appendix A we’ll show how one arrives to the solutions (67) and (72).
@@ Acknowledgements.
We are grateful to A. M. Kamchatnov for sending us the manuscript of his (to be published) excellent book ”Theory of Nonlinear Waves”, which motivated the continuation of our previous studies. We are very grateful to J. Cuevas-Maraver, M. Goldstein, A. Kamchatnov, E. Konig, B. Malomed, A. Novic-Cohen and V. Verbuk for the insightful discussions. We are also grateful to the organizers of Nor-Amberd School in Theoretical Physics, the scientific atmosphere of which stimulated the presented research. The paper was finalized during the Author’s participation in School and Workshop on Dynamical Systems, ICTP, Trieste. The Author is grateful to the Center for the hospitality.
Appendix A The ODE which doesn’t contain explicitly the independent variable
Equation (59) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $Q$ as the new independent variable and
${\cal I}=\frac{dQ}{d\tau}$ (74)
as the new dependent variable. In the new variables (59) takes the form of Abel equation of the second kind polyanin
$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=F(Q).$ (75)
@@ Appendix A The ODE which doesn’t contain explicitly the independent variable
$\lim_{Q\to Q_{1},Q_{2}}{\cal I}(Q)=0.$ (76)
@@ A.1 Integration in terms of elementary functions
$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\gamma(Q-Q_{1})(Q-Q_{2})(Q+Q_{3}).$ (77)
The boundary conditions (76) and the r.h.s. of (77) prompt the idea kogan4 ; kogan5 to look for ${\cal I}(Q)$ as the polynomial function proportional to $\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$. Because the r.h.s. of (77) is the third order polynomial of $V$, the multiplying factor should be a constant
${\cal I}(Q)=\frac{\psi}{Q_{1}-Q_{2}}\left(Q-Q_{1}\right)\left(Q-Q_{2}\right)$ (78)
@@
$\frac{T^{2}\psi^{2}}{\left(Q_{1}-Q_{2}\right)^{2}}\left(2Q-Q_{1}-Q_{2}\right)+\frac{\tau_{R}\psi}{Q_{1}-Q_{2}}=\gamma\left(Q+Q_{3}\right).$
@@
Hence the condition on the parameters of the problem, for (78) to be the solution of (77) is given by (71), and the constant multiplier in (78) should is given by (73). Substituting (78) into Eq. (74) we obtain the solution (72).
For $Q(V)$ given by (63) we obtain, instead of (77),
$T^{2}{\cal I}\frac{d{\cal I}}{dQ}+\tau_{R}{\cal I}=\beta(Q-Q_{1})(Q-Q_{2}).$ (80)
@@
If ${\cal I}$ is a polynomial function of $Q$, it should be a 3/2 order polynomial. And the boundary conditions (76) tell us that this polynomial should zero out at $Q=Q_{1,2}$.
Let us split in two one multiplier in the solution (78)
$\frac{Q-Q_{1}}{Q_{1}-Q_{2}}=\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}+1\right].$
This prompts us the idea to look up for the solution of (80) in the form kogan4 ; kogan5
${\cal I}(Q)=2\chi\left[\left(\frac{Q-Q_{2}}{Q_{1}-Q_{2}}\right)^{1/2}-1\right]\left(Q-Q_{2}\right).$ (82)
@@
Substituting the ansatz into (80) we see that it is in fact the solution of the equation, provided the parameters of the problem satisfy the relation (66) and $\chi$ is given by (68). Substituting thus obtained ${\cal I}(Q)$ into (74) we obtain (67)
A.2 Solitary wave solution of a generalized KdV equation
To additionally illustrate the method of integration used in this Appendix, let us recover in the framework of the method the solitary wave solution of a generalized KdV equation drazin
$\phi_{t}+(n+1)(n+2)\phi^{n}\phi_{x}+\phi_{xxx}=0.$ (83)
For a travelling wave solution
$\phi(t,x)=\phi(\tau),\hskip 28.45274pt\tau=t-x/U,$ (84)
the equation in partial derivatives (83) is reduced to the ODE
$U^{3}\frac{d\phi}{d\tau}-U^{2}(n+1)(n+2)\phi^{n}\frac{d\phi}{d\tau}-\frac{d^{3}\phi}{d\tau^{3}}=0.$ (85)
Equation (85) doesn’t contain explicitly the independent variable $\tau$. This prompts the idea to consider $\phi$ as the new independent variable and
$E=\frac{d\phi}{d\tau}$ (86)
as the new dependent variable. As the result, we can present the nonlinear third order ODE (85) as the linear second order ODE for $E^{2}$
$\frac{d^{2}E^{2}}{d\phi^{2}}=2U^{3}-2(n+1)(n+2)U^{2}\phi^{n}.$ (87)
A solitary wave corresponds to the boundary conditions
$\lim_{\tau\to\pm\infty}\phi(\tau)=0,$ (88)
which in the new variables are kogan5
$E^{2}(0)=0,\hskip 28.45274pt\left.\frac{dE^{2}}{d\phi}\right|_{\phi=0}=0.$ (89)
Solving (87) with these boundary conditions we obtain
$E^{2}=U^{3}\phi^{2}-2U^{2}\phi^{n+2}.$ (90)
Substituting thus obtained $E(\phi)$ into (86) and integrating we obtain the well known result drazin
$\phi^{n}=\frac{U}{2}\operatorname{sech}^{2}\left[\frac{n\sqrt{U}}{2}\left(x-Ut\right)\right].$ (91)
Appendix B Shock profile for the fully nonlinear transmission line
For the fully nonlinear transmission line it is convenient to present the problem in the form slightly different from that used everywhere above. We’ll consider the shock speed $U$ and the conditions before the shock as given. Thus the boundary conditions are
$\lim_{\tau\to-\infty}Q(\tau)=Q_{1};\hskip 14.22636pt\lim_{\tau\to-\infty}\Phi(\tau)$ $=\Phi_{1}$ (92)
and we can concretize (48) to
$\frac{1}{R_{L}}\frac{d\Phi}{d\tau}$ $=U(Q-Q_{1})-I+I(\Phi_{1}),$ (93a)
$R_{C}\frac{dQ}{d\tau}$ $=U(\Phi-\Phi_{1})-V+V(Q_{1}).$ (93b)
$V=U(\Phi-\Phi_{1})+V(Q_{1}).$ (94)
Substituting $V$ into (4a), solving the equation with respect to $Q$ and then substituting thus obtained $Q(\Phi)$ into (93a) we obtain the equation which can be easily integrated in quadratures. In the second case we equate the r.h.s. of (93a) to zero and obtain
$I=U(Q-Q_{1})+(\Phi_{1}).$ (95)
Substituting $I$ into (4b), solving the equation with respect to $\Phi$ and then substituting thus obtained $\Phi(Q)$ into (93b) we obtain the equation which can be integrated in quadratures. Thus we obtained Eqs. (48a) and (48b) of Ref. kogan3 .
If both $R_{C}>0$ and $R_{L}<\infty$, from (93) we may obtain the equation
$R_{L}R_{C}\frac{dQ}{d\Phi}=\frac{U(\Phi-\Phi_{1})-V(Q)+V(Q_{1})}{U(Q-Q_{1})-I(\Phi)+I(\Phi_{1})}.$ (96)
If we manage to solve (96), we can substitute the solution into one of the equations (93) and integrate the resulting equation in quadratures to obtain the profile of the shock wave. It is reassuring to realize that the conditions for the integrability in elementary functions of Eqs. (64) and (70) describing weak shocks in half-nonlinear transmission line (i.e. Eqs. (66) and (71)) are actually the (nonlinear) conditions on the product $R_{L}R_{C}$. Further analysis of (96) we postpone until later.